# Multiscale Modeling of Vacancy-Cluster Interactions and Solute Clustering Kinetics in Al-Mg-Zn Alloys

### v1

@@ Abstract
During heat treatment of age-hardenable alloys, growing solute clusters can strongly trap excess vacancies, limiting their mobility and significantly influencing precipitation kinetics. However, quantitative understanding of vacancy-cluster interactions remains limited due to the inherently multi-time- and multi-length-scale nature of vacancy-mediated diffusion. In this work, we develop an integrated computational framework combining lattice kinetic Monte Carlo (KMC) simulations, an atomistic absorbing Markov chain model, and mesoscale cluster dynamics (CD) to investigate these interactions and predict their effects on long-time clustering in Al-Mg-Zn alloys. The Markov chain model yields vacancy escape times from solute clusters and identifies a two-stage behavior of the vacancy-cluster binding energy: a size-dependent regime for small clusters (fewer than $\sim$100 atoms), followed by a saturation stage governed by vacancy formation energy differences between the cluster and matrix. These binding energies are used to estimate residual mobile vacancy concentrations in the Al matrix after quenching, which serve as critical inputs to CD simulations to predict the long-time cluster evolution kinetics during natural aging. Our results quantitatively demonstrate that cooling rate substantially affects natural aging kinetics, offering predictive insights for tailoring heat treatments in engineering alloys.
@@ Introduction
These complex interactions between vacancies and growing solute clusters in multicomponent alloys remain relatively underexplored by computational methods due to their chemical complexity and the inherently multi-time- and multi-length-scale nature of the underlying diffusion processes. Atomistic methods, such as on-lattice kinetic Monte Carlo (KMC) simulations, can accurately model vacancy migration kinetics using local-environment-dependent migration barriers informed by first-principles calculations [14, 15, 16]. However, these methods face inherent limitations in both spatial and temporal scales. Spatial constraints arise from the need to explicitly track individual atomic and vacancy jumps [17, 18], while temporal limitations stem from frequent low-barrier events that drastically reduce simulation efficiency [19, 20, 21]. Furthermore, conventional on-lattice KMC approaches struggle to capture the dynamic evolution of mobile vacancy concentrations [8, 9, 10], which is essential for predicting long-term diffusion and phase transformation behavior. To address these challenges, computational methods, such as phase-filed (PF) simulations [22, 23, 24], are needed to capture the feedback between evolving vacancy concentrations and solute clustering on the mesoscale. However, these mesoscale models require inputs such as mobility functions and initial cluster configurations, which ideally should be derived from atomistic simulations based on actual alloy compositions and heat treatment histories. Yet, this connection remains difficult due to the aforementioned limitations of atomistic methods. Consequently, a significant knowledge gap persists in linking atomic-scale vacancy-mediated diffusion with mesoscale phase transformation modeling.
Bridging the knowledge gap in vacancy-mediated diffusion processes is particularly critical for 7000 series Al-Mg-Zn alloys. These alloys can achieve high strengths (up to approximately 700 MPa ultimate tensile strength) after appropriate heat treatments [25]. However, their broader application, such as in the automotive industry,is limited because the rapid formation of solute clusters during natural aging significantly reduces ductility and formability [26, 27, 28]. Our recent studies [29, 30] revealed that Mg-Zn-rich solute clusters form rapidly during the quenching process, rather than only after reaching room temperature. Lattice KMC simulations show that these early-formed clusters strongly trap vacancies, leading to a significant reduction in the mobile vacancy concentration and potentially slowing subsequent precipitation kinetics [30]. This phenomenon, often referred to as “vacancy prisons” [31, 32], corresponds to deep energy basins on the potential energy landscape (PEL) that govern vacancy migration [11]. Importantly, however, vacancies trapped in these clusters are not completely immobile. The vacancy migration barriers within Mg-Zn-rich clusters are generally lower than in the surrounding Al matrix, allowing relatively fast vacancy diffusion within the cluster region and enabling a finite probability for vacancies to escape back into the matrix [30].
Although these qualitative descriptions provide useful insight into vacancy trapping, the overall effects on long-time solute clustering kinetics are still unclear due to the limitations of atomistic scale methods described above. A quantitative understanding of vacancy-cluster interactions remains limited. Zurob and Seyedrezai [33] proposed a model wherein the vacancy trapping effect is proportional to the cluster size (i.e., the number of solute atoms), which has been validated for small clusters in Al alloys [34, 35]. This relates to the growing spatial extent of the energy basin. In contrast, Soisson et al. [10] introduced a different approach for Cu clusters in the Fe-Cu system, emphasizing that the binding energy is governed by the vacancy formation energy difference in the Cu precipitate and the Fe matrix, which dictates the energetic depth of the trapping basin. While these models capture different aspects of the PEL, their application to multicomponent systems, such as Al-Mg-Zn, is hindered by the combinatorial complexity of cluster configurations and the complex interactions among different solute species. As a result, accurately modeling vacancy-cluster interactions in such systems remains a significant challenge, and conventional approaches (e.g., simple binding energy estimates) may not sufficiently account for this complexity [36, 11].
Fortunately, similar challenges involving low-energy basins have been extensively studied in the context of accelerating KMC simulations [37, 38, 39, 40, 41, 42, 43]. Conventional KMC approaches often become inefficient when systems are repeatedly trapped in states separated by small energy barriers, leading to prolonged residence in low-energy basins. To address this issue, several advanced methods have been developed. In particular, Chatterjee et al. [19], Puchala et al. [20], and Fichthorn et al. [21] introduced absorbing Markov chain frameworks that accelerate KMC simulations by efficiently treating transitions within such basins. Building upon these strategies, we apply an absorbing Markov chain model [20, 21] to systematically quantify the low-barrier, low-energy basins associated with Mg-Zn-rich solute clusters in Al-Mg-Zn alloys. In this approach, solute clusters are modeled as bidirected graphs, where nodes represent possible vacancy positions and edges denote transitions between nearest-neighbor lattice sites. By analyzing the transition probabilities and absorbing state behavior, we compute vacancy escape times from clusters and validate our results against direct KMC simulations. Additionally, we evaluate how effective vacancy binding energies vary with cluster size and assess their influence on the concentration of residual mobile vacancies in the Al matrix, a critical factor governing long-time diffusion kinetics after quenching. Our Markov chain model reveals a two-stage behavior in vacancy-cluster binding energies: an initial size-dependent regime for small clusters, consistent with the model proposed by Zurob and Seyedrezai [33], followed by a saturation regime governed by the vacancy formation energy difference between the cluster and matrix, in line with the findings of Soisson et al. [10]. Further details are provided in the Results section.
Finally, by leveraging the effective vacancy binding energies as functions of cluster size, which is quantified through the integrated KMC and absorbing Markov chain framework described above, we obtain essential input parameters for mesoscale simulations to capture long-time diffusion kinetics and associated diffusional phase transformations. cluster dynamics (CD) simulations provide a mesoscale modeling approach to describe the temporal evolution of atomic clusters (e.g., solute-vacancy clusters, precipitates, or defect aggregates) using population-based rate equations, without explicitly tracking individual atoms [44, 45, 46, 47, 48]. These rate equations incorporate the vacancy-cluster binding energy functions derived from our atomistic framework. Accordingly, we construct CD models to investigate long-time clustering kinetics, allowing mobile vacancy concentrations to evolve dynamically with the solute cluster configurations. The predicted results correspond well to experimentally relevant time and length scales. Our findings quantitatively demonstrate that cooling rate strongly influences natural aging kinetics, offering predictive insights for optimizing heat treatments in engineering alloys. Additionally, the predicted cluster distributions provide critical initial conditions for solute nucleation and early-stage growth in PF simulations of precipitation kinetics [22, 23, 24].
@@ Results
Figure 1: Time evolution of the average energy change per atom $E$ of two typical KMC quenching simulations[30]. The curve above is at a constant temperature of 300 K from the first step, representing an infinitely fast (instantaneous) cooling simulation from a solid solution. The curve below is from a water-quenching simulation. The change in the simulation temperature is denoted by the colors defined in the right color bar. Snapshots illustrate the initial and final structures from each simulation. The length of the cubic simulation supercells in the simulation snapshots is 12 nanometers.
@@
Another significant difference between the instantaneous cooling and water quenching cases in the above KMC simulations lies in the solute clustering kinetics at low temperatures ($\sim$300 K). In the instantaneous cooling case, the system’s energy continues to decrease at 300 K, indicating ongoing, albeit slow, vacancy-mediated solute diffusion. In contrast, energy evolution ceases entirely in the water quenching case, suggesting that vacancies become effectively trapped by pre-formed solute clusters.
@@
To systematically study the mechanism behind vacancy trapping in Al-Mg-Zn alloys, we first aimed to determine the structures and energetics of the clusters. Although density functional theory (DFT) calculations offer reliable energetic estimations, identifying the most stable structures of clusters with the given $n_{\text{Mg}}$ and $n_{\text{Zn}}$ in Al-Mg-Zn multicomponent alloys remains challenging because the possible configurations of solute clusters explode exponentially with the increase of $n_{\text{Mg}}$ and $n_{\text{Zn}}$ due to the change of atomistic structures. Hence, we employed surrogate models trained on first-principles data (see Supplementary Note 1) and applied a simulated annealing (SA) heuristic [49] adapted from the Metropolis-Hastings algorithm [50]. This method efficiently explores configuration space while minimizing the average formation enthalpy per solute atom, denoted as $\bar{H}_{\text{form}}$ (see Supplementary Note 2).
@@
Figure 2: Analyses of cluster local energetics obtained from SA. (a)-(c) An example cluster presenting the inner structures. The cluster is colored to show the spatial distribution of element type, the local vacancy binding energy of each site $E^{\text{site}}_{\text{bind}}$, and the local average vacancy migration barrier of each site $E^{\text{site}}_{\text{mig}}$. (d) Kernel density estimation plot for local vacancy migration barrier of each bond $E^{\text{bond}}_{\text{mig}}$ of this cluster based on their migration type. Here, NN shell denotes the first nearest neighbor shell of the cluster. The yellow region marks the vacancy migration event inside clusters. Light blue denotes the event vacancy jumping from the surface of the cluster to the nearest neighbor shell. Dark blue represents the event vacancy jumping out of the nearest neighbor shell. Dashed lines denote the averaged value for each type, and the black dotted line denotes the vacancy migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}$, of 0.58 eV. (e) Conceptual energy landscape of a ‘vacancy prison’ generated by a solute cluster. The dashed line connects local minima, while the curved line connects both minima and transition states.
@@
After the structures of the clusters were obtained, we conducted a comprehensive analysis. Specifically, we examined the local vacancy binding energy at each lattice site, $E^{\text{site}}_{\text{bind}}$, and the local migration barriers associated with each atomic bond, $E^{\text{bond}}_{\text{mig}}$. Figs. 2a-c depict an example cluster extracted from the SA simulations, sliced to show its internal structure. Fig. 2a highlights the atomic arrangement of Al (blue), Mg (green), and Zn (orange). This cluster exhibits a structure similar to $L1_{2}$, but a Zn/Mg ratio of approximately 2 due to its high surface-area-to-volume ratio. Fig. 2b presents the spatial distribution of the local binding energy of each site, $E_{\text{bind}}^{\text{site}}$, calculated from the following expression:
@@
where $E(\text{Vac})$ denotes the energy of the supercell containing a vacancy at a particular site in the cluster, and $E(\text{X})$ represents the energy of the same supercell with the atom X to fill up the previous local vacancy site. Atom X is the element originally present at that location, which can be Al, Mg, or Zn. Similarly, $E({\text{Al}}_{N-1}\text{Vac})$ and $E({\text{Al}}_{N-1}\text{X})$ correspond to the energies of a larger, pure Al supercell with a vacancy and with atom X occupying the vacancy site, respectively. This equation compares the formation energy difference of a vacancy-free cluster to the same cluster but containing a vacancy at a particular site. The value of $E_{\text{bind}}^{\text{site}}$ quantifies how strongly a vacancy binds at a particular site relative to the solute that occupies the position. Specifically, in a dilute Al matrix, $E_{\text{bind}}^{\text{site}}=0$, indicating no additional binding for the vacancy. In contrast, within a cluster, $E_{\text{bind}}^{\text{site}}$ typically exhibits a negative value, suggesting that vacancies are more likely to be bound or trapped within the cluster compared to the Al matrix. Fig. 2c similarly presents the spatial distribution of the local average migration barriers at each site, $E_{\text{mig}}^{\text{site}}$. This value represents the arithmetic average of the 12 vacancy migration barriers, $E_{\text{mig}}^{\text{bond}}$, associated with bonds connecting to its 12 first-nearest neighbors, assuming the site is initially occupied by a vacancy. The results highlight numerous sites within the cluster that exhibit low average migration barriers, with values around $\sim 0.3$ eV. Vacancies located at these sites have notably lower migration barriers compared to those at sites near the cluster shell, which present higher values similar to the vacancy migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}$, of $0.58$ eV.
To further investigate the variation in migration barriers, we classify the migration barriers for each bond, $E_{\text{mig}}^{\text{bond}}$, into three distinct groups, that is, migrations within clusters (yellow), transitions from the cluster surface to the nearest neighbor shell (light blue), and jumps beyond the nearest neighbor shell (dark blue), as depicted in Fig. 2d. The distributions of $E_{\text{mig}}^{\text{bond}}$ for each group are plotted. Consistent with observations in Fig. 2c, migration barriers are typically lower within the cluster. In contrast, vacancies jumping out of the nearest neighbor shell exhibit barriers primarily in the range of 0.54 to 0.62 eV, closely matching the vacancy migration barrier in pure Al. Ultimately, Fig. 2 illustrates that the cluster’s energetics create a low-energy basin in the PEL, as depicted in Fig. 2e. The basin presents mainly a short-range effect, and the cluster-matrix boundary can be interpreted as defining the edge of this energy basin, at which $E_{\text{bind}}^{\text{site}}\approx 0$ and $E_{\text{mig}}^{\text{bond}}\approx E_{\text{mig}}^{\text{Al}}$.
@@
Figure 3: Workflow for constructing an absorbing Markov chain model to analyze vacancy-cluster interactions and predict vacancy escape times. (a) An example solute cluster showing atomic arrangements and element distributions, with colors indicating Al (blue), Mg (green), and Zn (orange). (b) A graph representation of the cluster, including an additional layer of neighboring atoms. Nodes correspond to atoms/vacancies at lattice sites, and edges depict bonds between nearest neighbors. (c) An absorbing Markov chain (bidirected graph) derived from the directed graph, with nodes colored by local binding energies $E_{\text{bind}}^{\text{site}}$ and edges shaded based on migration barriers between connected nodes (sites) $E_{\text{mig}}^{\text{bond}}$. The outermost layer represents absorbing states where vacancies are considered escaped.
Therefore, we construct a vacancy escape model based on an absorbing Markov chain framework (see Methods section for concepts), based on the energetics obtained from the surrogate model (Figs. 2b-c). The workflow used to build the absorbing Markov chain model, predict vacancy escape times, and analyze vacancy-cluster interactions is illustrated in Fig. 3. As depicted in Fig. 3a, it starts with solute clusters. We obtained cluster structures directly from SA, although clusters can similarly be derived from KMC or other atomistic simulation methods. Then, the atomic structure is transformed into a graph representation (Fig. 3b). Each atom is represented as a node, and each atomic bond as an edge. To accurately describe the event vacancy jumping out of the nearest neighbor shell (Fig. 2d), an additional layer of neighboring atoms surrounding the cluster is included. Next, the graph is parameterized to become a bidirected graph, as shown in Fig. 3c. This transformation introduces directionality into the graph, labeling the information of vacancy migration paths. The outermost layer in this bidirected graph represents absorbing states, which are viewed as states in the Al matrix, where vacancies are considered to have escaped from the cluster region. The local binding energy, $E_{\text{bind}}^{\text{site}}$, is calculated for each node to characterize the thermodynamic stability of vacancies at specific lattice sites. Similarly, vacancy migration barriers between connected nodes, $E_{\text{mig}}^{\text{bond}}$, are also computed to parameterize transition probabilities between different sites.
@@
Here $\boldsymbol{T}$ is the transient matrix containing transition probabilities among all transient states, which is a square matrix of size $n_{\text{tr}}\times n_{\text{tr}}$, where $n_{\text{tr}}$ is the number of transient states and determined by sites in the cluster and the nearest neighbors. $\boldsymbol{I}$ is an identity matrix that has the same size as $\boldsymbol{T}$. $\vec{p_{0}}$ represents the initial probabilities of transient states with size $n_{\text{tr}}\times 1$, and $\vec{\tau}$ is a vector of one-step escape times for each site, which has the same size as $\vec{p_{0}}$.
@@
where $z$ is the coordination number ($z=12$ for an face-centered cubic (FCC) lattice), $E^{\text{bond}}_{\text{mig},ij}$ is the migration barrier from node (lattice site) $i$ to node (lattice site) $j$, corresponding to one of the 12 vacancy migration barriers from a lattice site to the nearest neighbor sites used to calculate previously defined $E_{\text{mig}}^{\text{bond}}$. To obtain the values of elements $p_{i}$ from $\vec{p_{0}}$, an equilibration assumption has been employed [20, 21, 42]. It assumes that each transient state will be visited numerous times since the energy basin cannot be escaped easily, eventually reaching a local equilibrium within the basin. Once this local equilibrium is established, the probability of occupying any given transient state $i$ follows a Boltzmann distribution:
$p_{i}=\frac{\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)}{\sum_{j}\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)},$ (4)

### v2

# Multiscale Modeling of Vacancy-Cluster Interactions and Solute Clustering Kinetics in Multicomponent Alloys

@@ Abstract
Prediction of solute clustering kinetics in aged multicomponent alloys requires a quantitative understanding of complex vacancy-cluster interactions across multiple scales. Here, we develop an integrated computational framework combining on-lattice kinetic Monte Carlo (KMC) simulations, absorbing Markov chain models, and mesoscale cluster dynamics (CD) to investigate these interactions in Al-Mg-Zn alloys. The Markov chain model yields vacancy escape times from solute clusters and identifies a two-stage behavior of the vacancy-cluster binding energy. These binding energies are used to estimate residual vacancy concentrations in the Al matrix after quenching, which serve as critical inputs to CD simulations to predict long-term cluster evolution kinetics during natural aging. Our results quantitatively demonstrate the significant impact of quench rate on natural aging kinetics. Results provide insights to guide alloy chemistry, quench rates, and aging time at finite temperature to control the evolution of solute clusters and eventual precipitates in aged multicomponent alloys.
@@ Introduction
Age-hardenable alloys, such as 6000 and 7000 series aluminum (Al) alloys, derive their mechanical strength from the controlled formation of solute clusters and precipitates through appropriate heat treatments that activate vacancy-mediated diffusion. Rapid quenching after solution treatment introduces a high concentration of residual vacancies, which significantly accelerates early-stage precipitation kinetics by enhancing the mobility of substitutional solute atoms [1]. At typical solutionizing temperatures ($\sim$500 C), the equilibrium vacancy concentration in Al is approximately $\sim 10^{-4}$, but it drops to around $\sim 10^{-11}$ at room temperature [2, 3]. Achieving this equilibrium requires extended timescales due to slow annihilation of vacancies at dislocations and grain boundaries. As a result, vacancy supersaturation can persist for hours (or even days) after quenching [4, 5]. Beyond their basic role in solute diffusion, vacancies interact strongly with solute atoms and clusters [6, 7, 8, 9, 10], dynamically altering the population of mobile vacancies. For instance, trace Sn additions in Al-Mg-Si alloys can effectively trap vacancies and suppress the formation of Mg-Si clusters during natural aging [11, 12]. Moreover, growing solute clusters themselves can bind vacancies, which reduces the available mobile vacancy concentration and slows further clustering kinetics [13, 14, 15, 16]. Therefore, a quantitative understanding of vacancy-cluster interactions and their impact on vacancy evolution is essential for accurately modeling and controlling diffusion-driven precipitation processes in age-hardenable alloys [17, 18], which ultimately determine key mechanical properties such as strength and ductility.
These complex interactions between vacancies and growing solute clusters in multicomponent alloys remain relatively underexplored by computational methods due to their chemical complexity and the inherently multi-time- and multi-length-scale nature of the underlying diffusion processes. Atomistic methods, such as on-lattice kinetic Monte Carlo (KMC) simulations, can accurately model vacancy migration kinetics using local-environment-dependent migration barriers informed by first-principles calculations [19, 20, 21]. However, these methods face inherent limitations in both spatial and temporal scales. Spatial constraints arise from the need to explicitly track individual atomic and vacancy jumps [22, 23], while temporal limitations stem from frequent low-barrier events that drastically reduce simulation efficiency [24, 25, 26]. Furthermore, conventional on-lattice KMC approaches struggle to capture the dynamic evolution of mobile vacancy concentrations [13, 14, 15], which is essential for predicting long-term diffusion and phase transformation behavior. To address these challenges, computational methods, such as phase-field (PF) simulations [27, 28, 29], are needed to capture the feedback between evolving vacancy concentrations and solute clustering on the mesoscale. However, these mesoscale models require inputs such as mobility functions and initial cluster configurations, which ideally should be derived from atomistic simulations based on actual alloy compositions and heat treatment histories. Yet, this connection remains difficult due to the aforementioned limitations of atomistic methods. Consequently, a significant knowledge gap persists in linking atomic-scale vacancy-mediated diffusion with mesoscale phase transformation modeling.
Bridging the knowledge gap in vacancy-mediated diffusion processes is particularly critical for 7000 series Al-Mg-Zn alloys. These alloys can achieve high strengths (up to approximately 700 MPa ultimate tensile strength) after appropriate heat treatments [30]. However, their broader application, such as in the automotive industry, is limited because the rapid formation of solute clusters during natural aging significantly reduces ductility and formability [31, 32, 33]. Our recent studies [34, 35] revealed that Mg-Zn-rich solute clusters form rapidly during the quenching process, rather than only after reaching room temperature. On-lattice KMC simulations show that these early-formed clusters strongly trap vacancies, leading to a significant reduction in the mobile vacancy concentration and potentially slowing subsequent precipitation kinetics [35]. This phenomenon, often referred to as “vacancy prisons” [36, 37], corresponds to deep energy basins on the potential energy landscape (PEL) that govern vacancy migration [16]. Importantly, however, vacancies trapped in these clusters are not completely immobile. Vacancy migration barriers within Mg-Zn-rich solute clusters are generally lower than in the surrounding Al matrix, allowing relatively fast vacancy diffusion within the cluster region and enabling a finite probability for vacancies to escape back into the matrix [35].
Although these qualitative descriptions provide useful insight into vacancy trapping, the overall effects on long-time solute clustering kinetics are still unclear due to the limitations of atomistic scale methods described above. A quantitative understanding of vacancy-cluster interactions remains limited. Zurob and Seyedrezai [38] proposed a model wherein the vacancy trapping effect is proportional to the cluster size (i.e., the number of solute atoms), which has been validated for small clusters in Al alloys [39, 40]. This relates to the growing spatial extent of the energy basin. In contrast, Soisson et al. [15] introduced a different approach for Cu clusters in the Fe-Cu system, emphasizing that the binding energy is governed by the difference in vacancy formation energies between the Cu precipitate and the Fe matrix, which dictates the energetic depth of the trapping basin. While these models capture different aspects of the PEL, their application to multicomponent systems, such as Al-Mg-Zn, is hindered by the combinatorial complexity of cluster configurations and the complex interactions among different solute species. As a result, accurately modeling vacancy-cluster interactions in such systems remains a significant challenge, and conventional approaches (e.g., simple binding energy estimates) may not sufficiently account for this complexity [41, 16].
Fortunately, similar challenges involving low-energy basins have been extensively studied in the context of accelerating KMC simulations [42, 43, 44, 45, 46, 47, 48, 49]. Conventional KMC approaches often become inefficient when systems are repeatedly trapped in states separated by small energy barriers, leading to prolonged residence in low-energy basins. To address this issue, several advanced methods have been developed. In particular, Chatterjee and Voter [24], Puchala et al. [25], and Fichthorn and Lin [26] introduced absorbing Markov chain frameworks that accelerate KMC simulations by efficiently treating transitions within such basins. Building upon these strategies, we apply an absorbing Markov chain model [25, 26] to systematically quantify the low-barrier, low-energy basins associated with Mg-Zn-rich solute clusters in Al-Mg-Zn alloys. In this approach, solute clusters are modeled as bidirected graphs, where nodes represent possible vacancy positions and edges denote transitions between nearest-neighbor lattice sites. By analyzing the transition probabilities and absorbing state behavior, we compute vacancy escape times from clusters and validate our results against direct KMC simulations. Additionally, we evaluate how effective vacancy binding energies vary with cluster size and assess their influence on the concentration of residual mobile vacancies in the Al matrix, a critical factor governing long-time diffusion kinetics after quenching. Our Markov chain model reveals a two-stage behavior in vacancy-cluster binding energies: an initial size-dependent regime for small clusters, consistent with the model proposed by Zurob and Seyedrezai [38], followed by a saturation regime governed by the vacancy formation energy difference between the cluster and matrix, in line with the findings of Soisson et al. [15]. Further details are provided in the Results section.
@@ Results
The distinct energy evolution profiles in Fig. 1 also lead to different solute cluster configurations, as shown in the inset figures visualized by OVITO[55]: the instantaneous cooling case results in a higher density of smaller clusters, whereas the water quenching case produces fewer but larger clusters. Another significant difference between these two cases lies in the solute clustering kinetics at low temperatures ($\sim$300 K). In the instantaneous cooling case, the system’s energy continues to decrease at 300 K, indicating ongoing, albeit slow, vacancy-mediated solute diffusion. In contrast, energy evolution ceases entirely in the water quenching case, suggesting that vacancies become effectively trapped by solute clusters formed during quenching [35].
[…]
@@
To reveal the mechanism behind vacancy trapping in Al-Mg-Zn alloys, we first aimed to determine the structures and energetics of the clusters. Although density functional theory (DFT) calculations offer reliable energetic estimations [56], identifying the most stable structures of clusters with the given numbers of Mg ($n_{\text{Mg}}$) and Zn ($n_{\text{Zn}}$) atoms in Al-Mg-Zn multicomponent alloys remains challenging because the possible configurations of solute clusters explode exponentially with the increase of $n_{\text{Mg}}$ and $n_{\text{Zn}}$. Hence, we employed surrogate models trained on first-principles data to compute the energy of a supercell with a specified lattice occupation configuration (see Supplementary Note 1 [56]). A simulated annealing (SA) heuristic [57], adapted from the Metropolis-Hastings algorithm [58], was then applied to generate clusters (see Supplementary Note 2). This method efficiently explores configuration space while minimizing the average formation enthalpy per solute atom.
@@
Figure 2: Analyses of local cluster energetics from SA simulations based on surrogate models trained on first-principles data [56, 35]. (a)-(c) An example solute cluster showing its internal structures: (a) spatial distribution of elemental species, (b) local vacancy binding energy at each site, $E^{\text{site}}_{\text{bind}}$, and (c) local average vacancy migration barrier at each site, $E^{\text{site}}_{\text{mig}}$. (d) Probability density plots obtained using kernel density estimation (KDE) for the local vacancy migration barrier associated with each first-nearest-neighbor bond, $E^{\text{bond}}_{\text{mig}}$, in this example cluster. Three distributions are shown, corresponding to different migration event types defined in the inset schematic: yellow indicates migration events within the cluster interior; light blue denotes vacancy jumps from the cluster surface to its first nearest-neighbor (NN) shell; dark blue represents vacancy jumps from the NN shell to the surrounding matrix. Dashed colored lines indicate average values for each event type, and the black dotted line marks the migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}=0.58$ eV. (e) Schematic illustration of a conceptual “vacancy prison” in the potential energy landscape (PEL) of vacancy migration around a solute cluster. The dashed line traces local energy minima, while the solid curved line connects both minima and transition states. Curved double-headed arrows represent vacancy migration.
@@
After the structures of the clusters were obtained, we examined the local vacancy binding energy at each lattice site, $E^{\text{site}}_{\text{bind}}$, and the local vacancy migration barriers associated with each atomic bond, $E^{\text{bond}}_{\text{mig}}$, between two nearest-neighbor sites. Figs. 2a-c depict an example cluster extracted from the SA simulations, sliced to show its internal structure. Fig. 2a highlights the atomic arrangement of Al (blue), Mg (green), and Zn (orange). This cluster exhibits a structure similar to the reported Guinier-Preston (GP)-I zones [59], where the number of Mg-Zn bonds is maximized, although it lacks strict periodicity. Fig. 2b presents the spatial distribution of $E_{\text{bind}}^{\text{site}}$ calculated from the following expression:
@@
where $E(\text{Vac})$ denotes the energy of the supercell containing a vacancy at a particular site in the cluster, and $E(\text{X})$ represents the energy of the same supercell with the atom X to fill up the previous local vacancy site. Atom X is the element originally present at that site, which can be Al, Mg, or Zn. Similarly, $E({\text{Al}}_{N-1}\text{Vac})$ and $E({\text{Al}}_{N-1}\text{X})$ correspond to the energies of a larger, pure Al supercell with a vacancy and with atom X occupying the vacancy site, respectively. This equation compares the formation energy difference of a vacancy-free cluster to the same cluster but containing a vacancy at a particular site. The value of $E_{\text{bind}}^{\text{site}}$ quantifies how strongly a vacancy binds at a particular site relative to the solute that occupies the position. Specifically, in a dilute Al matrix, $E_{\text{bind}}^{\text{site}}=0$, indicating no additional binding for the vacancy. In contrast, within a cluster, $E_{\text{bind}}^{\text{site}}$ typically exhibits a negative value, suggesting that vacancies are more likely to be bound or trapped within the cluster compared to the Al matrix.
Fig. 2c similarly presents the spatial distribution of the local average vacancy migration barriers at each site, $E_{\text{mig}}^{\text{site}}$. This value represents the arithmetic average of the 12 vacancy migration barriers, $E_{\text{mig}}^{\text{bond}}$, associated with atomic bonds connecting each site to its 12 first-nearest neighbors, assuming the site is initially occupied by a vacancy. The results highlight numerous sites within the cluster that exhibit low average migration barriers, with values around $\sim 0.3$ eV. These values of $E_{\text{mig}}^{\text{site}}$ are notably lower than those at sites near the cluster shell, which present higher values similar to the vacancy migration barrier in pure Al, $E_{\text{mig}}^{\text{Al}}$, of $0.58$ eV.
To further quantify the variations in migration barriers among different site types, Fig. 2d presents the probability distributions of $E_{\text{mig}}^{\text{bond}}$ for three categories of vacancy migration events: (i) migrations within the cluster (yellow), (ii) transitions from the cluster surface to the surrounding nearest-neighbor shell (light blue), and (iii) jumps beyond the nearest-neighbor shell (dark blue). Consistent with Fig. 2c, migration barriers are generally lower within the cluster interior. In contrast, vacancy jumps beyond the nearest-neighbor shell predominantly exhibit barriers in the range of 0.54 to 0.62 eV, closely aligning with the migration barrier in pure Al. Finally, Fig. 2e provides a conceptual illustration of the PEL of vacancy migration, representing the so-called “vacancy prison” created by a solute cluster. An energy basin is spatially localized at the cluster scale, with its boundary effectively defined by the cluster-matrix interface, where $E_{\text{bind}}^{\text{site}}\approx 0$ and $E_{\text{mig}}^{\text{bond}}\approx E_{\text{mig}}^{\text{Al}}$. The energy basin strongly binds vacancies, significantly reducing their mobility and effectively trapping them. Although trapped, vacancies within these "prisons" are not completely immobile. They can still move around within the cluster due to relatively low migration barriers. However, their probability of escaping back into the matrix is greatly reduced, leading to prolonged trapping.
@@
Figure 3: Workflow for constructing an absorbing Markov chain model to analyze vacancy-cluster interactions and predict vacancy escape times. (a) Atomic configuration of an example solute cluster, showing atomic arrangements and elemental distributions, with colors indicating Al (blue), Mg (green), and Zn (orange). (b) Graph representation of the cluster, including the cluster atoms and an additional layer of neighboring atoms. Nodes correspond to lattice sites occupied by atoms or vacancies, and edges represent bonds between nearest neighbors. (c) Absorbing Markov chain (bidirected graph) constructed from the graph in (b), where nodes are colored by local vacancy binding energies, $E_{\text{bind}}^{\text{site}}$, and edges are shaded according to the migration barriers between connected sites, $E_{\text{mig}}^{\text{bond}}$. The outermost layer denotes absorbing states, representing sites where vacancies are considered to have escaped from the cluster region.
We construct a numerical model to quantitatively describe the vacancy escape process from the “vacancy prison” illustrated in Fig. 2d based on an absorbing Markov chain framework (see Methods section for the related concepts), using the energetics obtained from the surrogate model (Figs. 2b-c) [56, 35]. The workflow used to build the absorbing Markov chain model, predict vacancy escape times, and analyze vacancy-cluster interaction energetics is illustrated in Fig. 3. As depicted in Fig. 3a, we first obtained solute cluster structures directly from SA, although clusters can similarly be derived from KMC or other atomistic simulation methods. Then, the atomic structure is transformed into a graph representation (Fig. 3b). Each atom is represented as a graph node, and each atomic bond as a graph edge. To accurately describe the event vacancy jumping out of the nearest neighbor shell (Fig. 2d), an additional layer of neighboring atoms surrounding the cluster is included. Next, the graph is parameterized to become a bidirected graph, as shown in Fig. 3c. This transformation introduces directionality into the graph, labeling the information of vacancy migration paths. The outermost layer in this bidirected graph represents absorbing states, which are viewed as states in the Al matrix, where vacancies are considered to have escaped from the cluster region. The local binding energy, $E_{\text{bind}}^{\text{site}}$, is calculated for each node using the surrogate model to characterize the thermodynamic stability of vacancies at specific lattice sites. Similarly, vacancy migration barriers between connected nodes, $E_{\text{mig}}^{\text{bond}}$, are also computed using the surrogate model to parameterize transition probabilities between different sites.
@@
Here $\boldsymbol{T}$ is the transient matrix containing transition probabilities among all transient states, which is a square matrix of size $n_{\text{tr}}\times n_{\text{tr}}$, where $n_{\text{tr}}$ is the number of transient states and determined by sites in the cluster and the nearest neighbors. $\boldsymbol{I}$ is an identity matrix that has the same size as $\boldsymbol{T}$. $\vec{p_{0}}$ is a vector to represent the initial probabilities of transient states with a size of $n_{\text{tr}}\times 1$, and $\vec{\tau}$ is a vector of one-step escape times for each site, which has the same size as $\vec{p_{0}}$.
@@
where $z$ is the coordination number ($z=12$ for an face-centered cubic (FCC) lattice), $E^{\text{bond}}_{\text{mig},ij}$ is the migration barrier from node $i$ to node $j$, corresponding to one of the 12 vacancy migration barriers from a lattice site to a nearest neighbor site. To obtain the values of elements $p_{i}$ in $\vec{p_{0}}$, an equilibration assumption has been employed [25, 26, 47]. It assumes that each transient state will be visited numerous times since the energy basin cannot be escaped easily, eventually reaching a local equilibrium within the basin. Once this local equilibrium is established, the probability of occupying any given transient state $i$ follows a Boltzmann distribution:
$p_{i}=\frac{\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)}{\sum_{j=1}^{n_{\text{tr}}}\exp\left({-\frac{E^{\text{site}}_{\text{bind},i}}{k_{B}T}}\right)},$ (4)